OpenStudy (anonymous):

Which point is 8 units from (–4, 7) on a coordinate plane? A. (–7, –4) B. (–4, –1) C. (–4, –7) D. (–4, 1)

OpenStudy (anonymous):

7-8=?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

kk

OpenStudy (welshfella):

use the formula for distance between 2 point for the 4 choices given d = sqrt [ (y2-y1)^2 + (x2 - x1)^2]

OpenStudy (welshfella):

eg for A it is sqrt[(7 - -4)^2 + (-4 --7)^2] = sqrt (11^2 + 3^2) = sqrt 130 which is not = 8 so not A

OpenStudy (welshfella):

Now try B